A man and his wife appear for an interview for two posts. The probability of husband’s selection is (1/7)and that of the wife’s selection is (1/5). Find the probability that only one of them will be selected.

Grade:Upto college level

2 Answers

SHAIK AASIF AHAMED
9 years ago
probability that only one of them is selected is given by
=(1/7)(4/5)+(1/5)(6/7)
=2/7.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty
SHAIK AASIF AHAMED
9 years ago
probability that only one of them is selected is given by
=(1/7)(4/5)+(1/5)(6/7)
=2/7.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty